i need help with 5-2e^(x+1)=1
Answer by RAY100(1637)

5-2e^(x+1) = 1,,,,,,subtract 5 both sides
.
-2e^(x+1) = 1-5 = -4,,,,,divide thru by (-2)
.
e^(x+1) = 2,,,,,,take ln both sides
.
(x+1) = ln 2 = .693,,,,,subtract 1 both sides
.
x= -.307,,,,,answer
.
.
check,,,,5 -2e^{(-.307)+1} =1
.
5-2e^{.693} =1
.
5-2{2) =1
.
5-4= 1,,,,,,,ok
.